An investor wants a real rate of return i1 of 6% per

year. If the expected annual inflation rate for the next

several years is 2.5%, what interest rate i should be

used in project analysis calculations?

Use : i= i'+f+(i'*f)
Answer #1

An investor will use nominal interest rate for an investment analysis

Nominal interest rate=Real interest rate+inflation rate+inflation rate*real interest rate

nominal interest rate=0.06+0.025+0.06(0.025)=0.06+0.025+0.0015=0.08650

Thus nominal interest rate=8.65%